# “I Can See Forever!”: Evaluating Real-time VideoLLMs for Assisting Individuals with Visual Impairments

### v1

@@ Abstract
The visually impaired population, especially the severely visually impaired, is currently large in scale, and daily activities pose significant challenges for them. Although many studies use large language and vision-language models to assist the blind, most focus on static content and fail to meet real-time perception needs in dynamic and complex environments, such as daily activities. To provide them with more effective intelligent assistance, it is imperative to incorporate advanced visual understanding technologies. Although real-time vision and speech interaction VideoLLMs demonstrate strong real-time visual understanding, no prior work has systematically evaluated their effectiveness in assisting visually impaired individuals. In this work, we conduct the first such evaluation. First, we construct a benchmark dataset (VisAssistDaily), covering three categories of assistive tasks for visually impaired individuals: Basic Skills, Home Life Tasks, and Social Life Tasks. The results show that GPT-4o achieves the highest task success rate. Next, we conduct a user study to evaluate the models in both closed-world and open-world scenarios, further exploring the practical challenges of applying VideoLLMs in assistive contexts. One key issue we identify is the difficulty current models face in perceiving potential hazards in dynamic environments. To address this, we build an environment-awareness dataset named SafeVid and introduce a polling mechanism that enables the model to proactively detect environmental risks. We hope this work provides valuable insights and inspiration for future research in this field.
Figure 1: Example of the experiment: a user with visual impairments navigating stairs with VideoLLMs.
@@ Introduction
Currently, both industry and academia have developed numerous devices and software aimed at providing support for individuals with severe visual impairments. Based on their functionality, these tools can be categorized into four classes: navigation aids [30, 21], image recognition applications [22, 9], daily life assistive products [15], and software-based tools [23]. While these technologies offer valuable support to visually impaired individuals, they share a common limitation: the inability to truly understand the semantics of the environment or contextualized tasks. The possession of powerful language understanding capabilities of Large Language Models (LLMs) and Vision Language Models (VLMs) offers a new solution to this problem. With the emergence of LLMs and VLMs such as GPT-4o, many researches have begun exploring their potential in assisting visually impaired individuals [37, 39, 17, 1]. However, most of them focus on understanding and describing static image scenes without fully addressing a more crucial capability in real-world scenarios, i.e., continuous perception. That is, how to utilize dynamic visual inputs such as videos to provide visually impaired users with a coherent and temporally consistent understanding and assistance. This gap limits the practical usability of assistive systems in complex and ever-changing environments.
The emergence of Video-based Large Language Models (VideoLLMs) [26] presents a promising opportunity to address these gaps. Trained on large-scale video-text pairs, VideoLLMs have demonstrated powerful video understanding capabilities. Models such as GPT-4o [16] and VITA-1.5 [11] exemplify this progress. These models enable real-time vision and speech interaction, laying the groundwork for their application in real-world assistive tasks for severely visually impaired Individuals. To evaluate whether state-of-the-art (SOTA) real-time VideoLLMs can play a substantive role in the daily life scenarios of severely visually impaired individuals, we conduct the first systematic empirical study.
@@
To the best of our knowledge, we present the first systematic evaluation of real-time VideoLLMs’ effectiveness in assisting visually impaired individuals (as shown in Figure 1). Firstly, based on previously established evaluation standards [27, 3] and insights from surveys conducted with severely visually impaired individuals, we categorize the evaluation dimensions into three main groups: (1) Basic Skills (including orientation skills, guided walking, independent walking, and cane techniques); (2) Home Life Tasks (such as housework, leisure, and recreation); and (3) Social Life Tasks (including road walking, transportation, and reaching destinations). For each task, we design corresponding test scenarios and task completion goals, resulting in the creation of a dataset named VisAssistDaily (Benchmark for Daily Assistance to the Severely Visually Impaired). We select three models capable of real-time vision and speech interaction for bilingual (English and Chinese) evaluation: GPT-4o [16], VITA-1.5 [11], and Zhipu [2]. We use four metrics: Task Success Rate, Prompt Cost, Response Latency, and Language Consistency. Experiments show that GPT-4o performs the best across all metrics, demonstrating a high task success rate and strong adaptability to both English and Chinese inputs. In contrast, VITA-1.5 records the lowest task success rate and performs poorly in language consistency, with instances of responding in Chinese to English prompts.
@@
In addition, we recruit visually impaired volunteers of varying ages, genders, and degrees of vision loss to participate in the evaluation. The volunteers participate in both closed-world and open-world tests and provide feedback on the VideoLLMs. The results show that users are generally satisfied with the response speed. Users are generally satisfied with the response speed, ease of use, and pricing. However, they express concerns about the accuracy in complex environments, such as stair recognition. Based on the evaluations, we identify key challenges and future directions in this domain. To address the “Proactive Perception” challenge, we propose a simple yet effective solution by constructing a novel environment-awareness dataset, SafeVid, which integrates a diverse set of real-world hazard scenarios, such as approaching obstacles and contact with dangerous objects. Then, we fine-tune the VITA-1.5 on this dataset and implement a polling-based query mechanism, enabling the model to proactively alert the environmental risks. The fine-tuned model achieves an accuracy of 62.24%, showing its effectiveness in real-world applications.
In conclusion, we make the following contributions:
@@
  • We conduct the first systematic evaluation of the practical effectiveness of VideoLLMs in assisting severely visually impaired individuals with daily life tasks. To support standardized testing, we design a benchmark called VisAssistDaily, which measures model performance in multilingual scenarios in terms of task completion and response quality.
  • We design a user study, inviting different degrees of visually impaired volunteers from diverse backgrounds to test VideoLLMs. By collecting their feedback during actual interactions, we further validate the strengths and limitations of these models in practical use cases.
@@
  • Through in-depth analysis of the experimental process and user feedback, we identify key challenges that VideoLLMs currently face in this domain. To address the lack of proactive prompting in VideoLLMs, we propose SafeVid, a dataset for training models to alert users of potential risks.
Figure 2: Overview of dataset and examples.
@@
With the emergence of LLMs and VLMs, many researchers leverage their powerful understanding capabilities to provide better support for visually impaired individuals. Kuzdeuov et al. [17] develop a mobile application specifically designed for visually impaired users, allowing them to interact with ChatGPT through natural speech. This enables easier access to information and task completion, helping users live and learn more independently. Zhang et al. [37] designed a novel interactive system for the visually impaired that combines image semantic segmentation, language models, and haptic feedback. This system allows users to “see” objects and semantic information in images through voice and touch. Zhao et al. [39] proposed a task framework named VIALM (Visually Impaired Assistance with Large Models), focusing on the application of large models in visual impairments assistance. However, real-world environments are dynamic. Relying solely on static image understanding is insufficient to meet the needs of visually impaired individuals for continuous perception in real-life situations. There is still a lack of in-depth exploration of continuous visual inputs such as video, which is the central focus of this study.
@@
This paper designs and constructs two datasets. The first is an evaluation benchmark, VisAssistDaily, which is used to assess the ability of VideoLLMs to assist severely visually impaired individuals in completing daily activities. The second dataset is a video-text paired dataset, SafeVid, intended to support the model in proactively perceiving and alerting users to potential hazards in the environment.
@@
It focuses on evaluating whether VideoLLMs can assist the visually impaired individuals in performing basic daily mobility skills. The overall process, illustrated in the Figure 3, includes four tasks encompassing Orientation Skills, Guided Walking, Independent Walking, and Cane Techniques.
@@
These tasks evaluate the model’s ability to support individuals with visual impairments in completing household tasks and engaging in reading activities within a home environment, aiming to enhance their daily independence. As shown in the Figure 4, the overall process is divided into two parts:
@@
This section simulates mobility tasks for visually impaired individuals during social activities, testing the model’s ability to support path recognition, transportation assistance, and destination guidance in real-world scenarios. As illustrated in the Figure 5 , the evaluation consists of three parts:
@@
To enable proactive risk alert functionality for VideoLLMs, we collect 1,204 high-definition 1080p video clips from real-world environments and annotate them based on the severity of potential environmental hazards present in each scene as shown in  Figure 6.
As shown in Figure 2, the collected videos cover typical locations visited by visually impaired individuals, such as bus stops, sidewalks, banks, and restaurants. Based on these data, we establish a risk-level classification system that categorizes each video according to the potential severity of future hazards. The system includes the following four levels:
@@
Figure 6: SafeVid’s dataset construction example.
Table 1: Model performance on the VisAssistDaily benchmark across assistive tasks in three categories and two languages (EN and CH).
Category Task GPT-4o Zhipu VITA-1.5
@@
VideoLLMs Evaluation. We evaluate the assistive potential of three SOTA VideoLLMs, comprising both production-level and open-source systems, in supporting severely visually impaired individuals. (1) GPT-4o executes through the ChatGPT iOS application (v1.2025.043), (2) Zhipu is used through its iOS client (v2.7.0), (3) VITA-1.5 can be found in its Github  [^1] https://github.com/VITA-MLLM/VITA..
Input Devices. In this paper, we use an “iPhone 13 Pro Max in iOS 18.3.1 system” 222 https://www.apple.com/by/iphone-13-pro/specs/. as the primary testing device. The app/website with the VideoLLMs are executed on this device, and first-person video data is captured using its built-in screen recording function. To obtain a more experimental record, we adopt a dual-device recording setup by introducing a “DJI Action 4”  [^3] https://www.dji.com/cn/osmo-action-4. camera mounted on a stabilizing rig to simultaneously capture video of a third-person perspective.
@@
Table 3: Summary of results from the user study, including age, visual status, task success rates in closed-world (TSR (C)) and open-world (TSR (O)) scenarios, and satisfaction scores (Sat.). Visual Status abbreviations: B = Blind, SVI = Severely Visually Impaired, S = Sighted, youth = since youth, later = later in life, bf = blindfolded.
ID Age Visual Status TSR (C) TSR (O) Sat.
U1 90 B (youth) 60.00% 75.00% 7
U2 35 B (youth) 100.00% 66.67% 8
U3 30 SVI (child) 75.00% 50.00% 9
U4 21 SVI (child) 80.00% 83.33% 7
U5 27 B (later) 80.00% 100.00% 6
U6 24 S (bf) 100.00% 100.00% 7
U7 22 S (bf) 100.00% 100.00% 8
U8 24 S (bf) 100.00% 80.00% 8

### v2

@@ Abstract
The visually impaired population faces significant challenges in daily activities. While prior works employ vision language models for assistance, most focus on static content and cannot address real-time perception needs in complex environments. Recent VideoLLMs enable real-time vision and speech interaction, offering promising potential for assistive tasks. In this work, we conduct the first study evaluating their effectiveness in supporting daily life for visually impaired individuals. We first conducted a user survey with visually impaired participants to design the benchmark VisAssistDaily for daily life evaluation. Using VisAssistDaily, we evaluate popular VideoLLMs and find GPT-4o achieves the highest task success rate. We further conduct a user study to reveal concerns about hazard perception. To address this, we propose SafeVid, an environment-awareness dataset, and fine-tune VITA-1.5, improving risk recognition accuracy from 25.00% to 76.00%. We hope this work provides valuable insights and inspiration for future research in this field.
Figure 1: Overview of the experiment setup and interaction flow with examples of VideoLLM.
@@ Introduction
Prior works have developed numerous devices and software aimed at providing support for individuals with severe visual impairments [36]. Based on their functionality, these tools can be categorized into four classes: navigation aids [38, 25, 19, 6, 18], image recognition applications [26, 10, 28], daily life assistive products [16], and software-based tools [29]. While these technologies offer valuable support to visually impaired individuals, they share a common limitation: the inability to truly understand the semantics of the environment or contextualized tasks. The possession of powerful language understanding capabilities of Large Language Models (LLMs) and Vision Language Models (VLMs) offers a new solution to this problem. With the emergence of LLMs and VLMs, many researches have begun exploring their potential in assisting visually impaired individuals [46, 48, 20, 1, 41]. However, most of them focus on understanding and describing static image scenes without fully addressing a more crucial capability in real-world scenarios, i.e., continuous perception. In practice, visually impaired individuals receive information through dynamic and ever-changing inputs, such as ongoing video streams. Therefore, it is essential for models to handle sequential visual inputs and deliver assistance that is both temporally consistent and practically useful. This gap limits the practical usability of assistive systems in complex and ever-changing environments.
The emergence of Video-based Large Language Models (VideoLLMs) [33] presents a promising opportunity to address these gaps. Trained on large-scale video-text pairs, VideoLLMs have demonstrated powerful video understanding capabilities. Currently, VideoLLMs can be broadly categorized into offline and real-time modes. The latter, exemplified by models such as GPT-4o [17] and VITA-1.5 [12] exemplify this progress. These models enable real-time vision and speech interaction, laying the groundwork for their application in real-world assistive tasks for visually impaired individuals. [^1] In this paper, all mentions of VideoLLMs refer to real-time VideoLLMs. To evaluate whether state-of-the-art (SOTA) real-time VideoLLMs can play a substantive role in the daily life scenarios of visually impaired individuals, we conduct the first systematic empirical study.
@@
To the best of our knowledge, we present the first systematic evaluation of real-time VideoLLMs’ effectiveness in assisting visually impaired individuals (as shown in Figure˜1). Firstly, based on established evaluation standards [34, 3] and our user survey with visually impaired individuals (including a questionnaire and short semi-structured interviews), we categorize the evaluation dimensions into three groups: (1) Basic Skills (including orientation skills, guided walking, independent walking, and cane techniques); (2) Home Life Tasks (such as housework, leisure, and recreation); and (3) Social Life Tasks (including road walking, transportation, and reaching destinations). For each task, we design corresponding test scenarios and task completion goals, resulting in the creation of a dataset named VisAssistDaily (Benchmark for Daily Assistance to the Visually Impaired). We select three VideoLLMs capable of real-time vision and speech interaction for bilingual (English and Chinese) evaluation: GPT-4o [17], VITA-1.5 [12], and Zhipu [2]. We consider four metrics in our evaluation: Task Success Rate, Prompt Cost, Response Latency, and Language Consistency. Experiments show that GPT-4o performs the best across all metrics, demonstrating a high task success rate and strong adaptability to both English and Chinese inputs. In contrast, VITA-1.5 records the lowest task success rate and performs poorly in language consistency, with instances of responding in Chinese to English prompts.
@@
In addition, we recruit visually impaired volunteers of different ages, genders, and levels of vision loss to participate in our user study. Guided by the three evaluation dimensions defined earlier, the volunteers complete both indoor and outdoor tasks and provide structured feedback on the performance of VideoLLMs. The results indicate that users are generally satisfied with the response speed, ease of use, and affordability, but express concerns about accuracy in complex environments, particularly in stair recognition. Based on previous evaluations, we identify key challenges and future directions in this domain. To address the “Proactive Perception” challenge, we construct a novel environment-awareness dataset, SafeVid, which incorporates diverse real-world hazard scenarios such as approaching obstacles and contact with dangerous objects. We then fine-tune VITA-1.5 on SafeVid for the environmental risk recognition task, improving accuracy from 25.00% to 76.00% and demonstrating the effectiveness of our approach in real-world applications. In conclusion, we make the following contributions:
@@
  • We conduct the first systematic evaluation of the practical effectiveness of VideoLLMs in assisting visually impaired individuals with daily life. To support this, we carry out a user survey with visually impaired individuals and use the insights to design a benchmark, VisAssistDaily, which measures VideoLLMs’ performances in multilingual scenarios in terms of task completion and response quality.
  • We conduct a user study with visually impaired participants of varying backgrounds and degrees of vision loss. Through their real-world interactions with VideoLLMs, we collect authentic feedback that reveals both the strengths and limitations of these models in practical assistive scenarios.
@@
  • Through in-depth analysis of the experimental process and user feedback, we identify key challenges that VideoLLMs currently face in this domain. To address the most important challenge, lack of proactive prompting in VideoLLMs, we propose a dataset SafeVid, and then fine-tuned VITA-1.5 on the environmental risk recognition task, improving accuracy from 25.00% to 76.00%.
Figure 2: Overview of VisAssistDaily and SafeVid.
@@
With the emergence of LLMs and VLMs, many researchers leverage their powerful understanding capabilities to provide better support for visually impaired individuals. Kuzdeuov et al. [20] develop a mobile application specifically designed for visually impaired users, allowing them to interact with ChatGPT through natural speech. This enables easier access to information and task completion, helping users live and learn more independently. Zhang et al. [46] design a novel interactive system for the visually impaired that combines image semantic segmentation, language models, and haptic feedback. This system allows users to “see” objects and semantic information in images through voice and touch. Zhao et al. [48] propose a task framework named VIALM (Visually Impaired Assistance with Large Models), focusing on the application of large models in visual impairments assistance. However, real-world environments are dynamic. Relying solely on static image understanding is insufficient to meet the needs of visually impaired individuals for continuous perception in real-life situations. There is still a lack of in-depth exploration of continuous visual inputs such as video, which is the central focus of this study.
Table 1: Questionnaire on mobility, household, and social tasks. Symbols: ◐ denotes single-choice questions, ● denotes multiple-choice questions.
Domain Type Question Options
Basic Skills How difficult do you find it to orient yourself in unfamiliar environments? a) Very difficult  b) Mostly difficult  c) Moderate  d) Slightly difficult  e) Not difficult at all
Do you feel comfortable walking independently in familiar environments? a) Very dissatisfied  b) Dissatisfied  c) Neutral  d) Satisfied  e) Very satisfied
What problems do you commonly face when walking with a cane? a) Incomplete/invisible accessible paths  b) Too many obstacles (cars, objects)  c) Difficulty judging terrain changes (steps, slopes)  d) Lack of cane training  e) Others
Home Life Tasks What household tasks do you most often need help with? a) Cleaning  b) Organizing/storing items  c) Cooking/preparing food  d) Washing clothes  e) Others
What are your main means of leisure and entertainment? a) Listening to radio/audiobooks  b) Using smartphone/computer apps (screen reader, voice)  c) Chatting with family/friends  d) Community activities  e) Others
If given more training/support, in which areas would you like to improve your independence? a) Household management  b) Using tech products (phone, computer, smart speaker)  c) Time management/planning  d) Leisure/recreation skills  e) Others
Social Life Tasks What social barriers do you most commonly encounter? a) Difficult to reach venues  b) Lack of accessible facilities  c) Inaccessible info (no voice prompts)  d) Communication barriers  e) Lack of companionship  f) Others
How difficult is walking during social activities for you? a) Very difficult  b) Mostly difficult  c) Moderate  d) Slightly difficult  e) Not difficult at all
What are the main difficulties you face when using public transportation? a) Difficult to find station/entrance  b) Difficult to get transfer/arrival info  c) Inconvenient boarding/alighting  d) Lack of support from staff  e) Others
How would you assess your mobility in daily social life? a) Adequate, meets my needs  b) Fair, with room for improvement  c) Inadequate, often restricted
(a) Basic skills.
(b) Home Life Tasks.
(c) Social Life Tasks.
Figure 3: Illustrations for questionnaire items.
User Survey
To ensure that our evaluation of VideoLLMs accurately reflects the core needs of the individual with visual impairments, we design and conduct a questionnaire together with short semi-structured interviews. Participants are directly involved in the construction of the evaluation benchmark. The questionnaire covers three main themes, including daily mobility, home life activities, and social participation (as shown Table˜1), which grounds the design of VisAssistDaily in user needs and real-world challenges.
Set Up
Participants. We recruit a total of 50 adults with visual impairments, including 24 men (48%), 23 women (46%), and 3 non-binary or other individuals (6%). In terms of visual status, 24 participants are blind (48%) and 26 are low-vision (52%). All participants provide informed consent after being fully informed of the study purpose and data use.
Tools and Procedure. The questionnaire items are grouped into basic skills, household tasks, and social tasks, using single-choice or multiple-choice formats. Optional semi-structured interviews follow the questionnaire, focusing on recent challenges and expectations for AI assistance. The questionnaire is distributed to real blind communities and online platforms for blind communities, such as Reddit. In addition, two open-ended questions are analyzed through inductive thematic analysis, and word clouds are created as complementary visualizations (as shown Figure˜4 for more details). Synonyms are merged during processing, for example, “AI assistant”, “voice assistant”, and “smart life assistant” are combined into “AI/voice assistant.”
Results
Basic Skills. In the area of basic skills, participants report significant difficulties as shown in Figure˜3(a).
  1. 1.
    Orientation ability. In unfamiliar environments, most participants report at least moderate difficulty: 6 “very difficult”, 15 “difficult”, and 19 “moderate”, while 10 report “mostly no difficulty”. This underscores the need for VideoLLMs to recognize environmental features and provide reliable directional guidance.
  2. 2.
    Independent walking. In familiar environments, most respondents express satisfaction with their ability to walk independently: 21 “satisfied” and 12 “very satisfied”, with 9 “neutral” and 8 “dissatisfied”. Interview responses suggest this confidence depends heavily on environmental stability and drops quickly when obstacles or changes occur.
  3. 3.
    Challenges in cane walking. The most frequent difficulties selected are “many obstacles” (38) and “unclear access path” (28), followed by “undetected surface changes” (26) , “lack of cane training” (10), and “other” (2).
These findings highlight the importance of obstacle recognition, path guidance, and technique support in our benchmark tasks.
Household Life Tasks. In the area of household tasks, participants report significant difficulties as shown in Figure˜3(b).
  1. 1.
    Daily household needs. Cooking or meal preparation is selected 41 times, organizing or storing items 24 times, doing laundry 11 times, cleaning 9 times, and other activities 2 times. The results indicate that the main challenges in household life center on meal preparation and item management.
  2. 2.
    Leisure and entertainment activities. Using mobile or computer applications is selected 38 times, listening to audio content 31 times, chatting with family or friends 26 times, and participating in community activities 20 times. These results suggest that mobile devices have become the primary entry point for information access and leisure.
  3. 3.
    Desired skills to improve. The use of digital devices is selected 42 times, household management 24 times, time management or planning 18 times, leisure skills 13 times, and other activities 1 time.
These findings highlight the importance of household item management, mobile-based interaction, and digital device usage in our benchmark tasks.
Social Life Tasks. In the area of social life tasks, participants report significant difficulties as shown in Figure˜3(c).
  1. 1.
    Common social barriers. Reported frequencies are: travel difficulties 27, insufficient accessibility of venues 27, difficulties in accessing information 32, lack of companionship support 23, communication barriers 6, and other barriers 1. Overall, information accessibility and the lack of accessible facilities in public spaces emerge as the most prominent issues, while the results also reflect limited companionship support and social resources.
  2. 2.
    Walking difficulties in social activities. Very difficult 3, difficult 13, moderate 22, mostly no difficulty 11, no difficulty 1. Most participants report moderate or lighter levels of difficulty, suggesting that environmental conditions often play a more decisive role than personal walking ability.
  3. 3.
    Main difficulties in public transportation. Locating stops or entrances 22, transfer and arrival information 31, getting on or off vehicles 34, support from drivers or staff 14. The results indicate that the main obstacles in travel are concentrated in station recognition, access to transportation information, and the process of boarding and alighting.
  4. 4.
    Self-assessment of mobility in daily social life. Sufficient 6, average 29, insufficient 15. The majority of participants rate themselves as “average”, though a considerable proportion still consider their ability insufficient.
These findings highlight the importance of information access, accessible venues, and public transport support in our benchmark tasks.
Open-ended Results.
  1. 1.
    Q1: Areas to improve beyond mobility (n=33). The most frequently mentioned topics are reading and education accessibility (6/33) and software or digital accessibility (6/33), followed by employment and workplace support (4/33) and voice guidance in public or commercial venues (4/33). Other mentions include home and everyday product accessibility (3/33), signage and information presentation (3/33), accessibility infrastructure (2/33), indoor navigation and obstacle alerts (2/33), and several single mentions (1/33 each): spoken medication instructions, ride-hailing driver training, simplified operation modes, volunteer companionship, accessible recreational activities, and skills training (as shown Figure˜4(a)).
    Overall, the responses converge on two clusters: information and content accessibility (reading, software, signage, simplified modes) and contextual guidance for participation (voice guidance, workplace support). At the same time, the long tail of single mentions reveals diverse and highly individualized needs.
  2. 2.
    Q2: Expectations and concerns for future assistive technologies (n=23). Positive expectations highlight wide adoption and smarter systems (5/23) and AI or voice assistants (4/23), with single mentions of human-in-the-loop remote assistance, remote navigation, home-care tools, greater social acceptance, better environment recognition, stronger accessibility infrastructure, richer audio cue systems, and accessible workplaces. Concerns mainly focus on cost and affordability (3/23), privacy and data security (2/23), and usability or learning burden (2/23), with isolated mentions of latency and safety risks (1/23) and the pace of updates and adaptation (1/23)(as shown Figure˜4(b)).
    Users envision assistive systems that are both intelligent and ubiquitous, but their primary worries concern affordability, trust in data practices, and the ease of learning. This combination suggests a strong demand for mainstream, low-barrier, and privacy-conscious design.
(a) Word cloud results for Q1: areas to improve beyond mobility.
(b) Word cloud results for Q2: expectations and concerns for future assistive technologies.
Figure 4: Word clouds for the user survey’s open-ended results.
Table 2: Basic Skills (Tool: Cane).
         Image Task Goal Instruction
Location
& Tool
  [Uncaptioned image] Orientation Skills (a) Ask environment (b) Describe obstacles (c) Determine facing direction Evaluates the model’s ability to help users determine direction in complex environments. (a) I am a blind person. Please describe the environment in front of me. (b) Could you please tell me if there are any obstacles in front of me? (c) Could you tell me which direction I’m facing to? Outdoor / Yes
[Uncaptioned image] Guided Walking (a) Ask staircase details (b) Guide ascent (c) Go upstairs (d) Make it upstairs Tests if the model can safely guide stair climbing using real-time cane feedback. (a) I am a blind person. Please tell me what is in front of me. (b) Please guide me step by step as I climb the stairs. (c) I am starting to go up the stairs now. Please help me to do it. (d) I have reached the top of the stairs. Can you describe the area ahead? Outdoor / Yes
[Uncaptioned image] Independent Walking (a) Ask room layout (b) Identify entry conditions (c) Locate lost object Assesses whether the model can direct a user into a room without tools. (a) I am a blind person. Can you describe the room in front of me? (b) What should I do if I would like to go into this room? (c) I have lost my cane in the room. Can you help me to find it? Indoor / No
[Uncaptioned image] Cane Techniques (a) Navigate along wall (b) Retrieve cane (c) Guide to seating Evaluates the model’s ability to recognize and guide proper cane use. (a) I am a blind person. Please guide me to follow the wall. (b) I need to find and pick up my cane, can you help me to find it? (c) I have found my cane, please guide me to the nearest seat. Indoor / No
Table 3: Home Life Tasks (Tool: Cane).
         Image Task Goal Instruction
Location
& Tool
  [Uncaptioned image] Housework (a) Find object ahead (b) Determine position (c) Identify the property Tests the model’s ability to identify object locations and attributes at home. (a) I am a blind person. Please help me identify what objects are in front of me. (b) Can you tell me exactly where this object is located relative to me? (c) What kind of object is it? Indoor / No
[Uncaptioned image] Leisure and Recreation (a) Identify the books (b) Turn the pages (c) Read the table of contents Assesses the model’s ability to help access printed books through content recognition. (a) I am a blind person. Can you help me identify which book is in front of me? (b) Please guide me on how to turn the pages of this book. (c) Can you tell me how to find the page of the table of contents? Indoor / No
Table 4: Social Life Tasks (Tool: Cane).
         Image Task Goal Instruction
Location
& Tool
  [Uncaptioned image] Road Walking (a) Describe the environment (b) Walk on blind lane (c) Navigate by tapping a cane Evaluates the model’s ability to support safe navigation on roads using tactile paths. (a) I am a blind person. Please describe the surroundings around me. (b) I am on the blind lane, please guide me to walk safely along it. (c) As I tap my cane, can you help me figure out what’s around me and guide me forward? Outdoor / Yes
[Uncaptioned image] Transport (a) Ask location of the car (b) Find the door (c) Open the door (d) Get on the car Tests the model’s ability to assist with vehicle identification and boarding. (a) I am a blind person. Can you tell me where the car is parked? (b) Please guide me to locate the car door. (c) Now I reach it, please help me open the door. (d) Please guide me to get into the car. Outdoor / Yes
[Uncaptioned image] Reaching Destination (a) Recognize environment (b) Make sure the location of chair (c) Judge the chair’s position (d) Sit down Assesses whether the model can guide the user to complete final actions upon arrival. (a) I am a blind person. Can you help me recognize where I’ve arrived? (b) Is there a chair nearby? Please confirm its location. (c) Please help me to find the exact position of the chair. (d) Please guide me to sit down on the chair. Indoor / Yes
@@ Results
In this paper, we construct two datasets. The first is an evaluation benchmark, VisAssistDaily, which is used to assess the ability of VideoLLMs to assist severely visually impaired individuals in completing daily activities. The second dataset is a video-text paired dataset, SafeVid, intended to support the model in proactively perceiving and alerting users to potential hazards in the environment.
@@
It focuses on evaluating whether VideoLLMs can assist the visually impaired individuals in performing basic daily mobility skills. The overall process, illustrated in the Section˜3.2, includes four tasks encompassing Orientation Skills, Guided Walking, Independent Walking, and Cane Techniques.
@@
These tasks evaluate the model’s ability to support individuals with visual impairments in completing household tasks and engaging in reading activities within a home environment, aiming to enhance their daily independence. As shown in the Section˜3.2, the overall process is divided into two parts:
@@
This section simulates mobility tasks for visually impaired individuals during social activities, testing the model’s ability to support path recognition, transportation assistance, and destination guidance in real-world scenarios. As illustrated in the Section˜3.2 , the evaluation consists of three parts:
@@
To enable proactive risk alert functionality for VideoLLMs, we collect 2,205 high-definition 1080p video clips from real-world environments and annotate them based on the severity of potential environmental hazards present in each scene, as shown in Figure˜5.
Figure 5: Example of SafeVid dataset.
The collected videos cover typical locations visited by visually impaired individuals, such as bus stops, sidewalks, markets, and restaurants (See Figure˜2 for more details). Based on these data, we establish a risk-level classification system that categorizes each video according to the potential severity of future hazards. The system includes the following four levels:
@@
Table 5: Model performance on the VisAssistDaily benchmark across assistive tasks in three categories and two languages (EN: English and CH: Chinese).
Category Task GPT-4o Zhipu VITA‑1.5
@@
VideoLLMs Evaluation. We evaluate the assistive potential of three SOTA VideoLLMs, comprising both production-level and open-source systems, in supporting severely visually impaired individuals: (1) GPT-4o executes through the ChatGPT iOS application (v1.2025.043); (2) Zhipu is used through its iOS client (v2.7.0); (3) VITA-1.5 can be found in its Github.  [^2] https://github.com/VITA-MLLM/VITA.
Input Devices. In this paper, we use an “iPhone 13 Pro Max in iOS 18.3.1 system”  [^3] https://www.apple.com/by/iphone-13-pro/specs/. as the primary testing device. The app/website with the VideoLLMs are executed on this device, and first-person video data is captured using its built-in screen recording function. To obtain a more experimental record, we adopt a dual-device recording setup by introducing a “DJI Action 4”  [^4] https://www.dji.com/cn/osmo-action-4. camera mounted on a stabilizing rig to simultaneously capture video of a third-person perspective.
@@
Table 6: Participant demographics, including age, visual status, task success rates in closed-world (TSR (C)) and open-world (TSR (O)) scenarios, and satisfaction scores. Visual status abbreviations: B = Blind, SVI = Severely Visually Impaired, S = Sighted; onset abbreviations: Youth = Since Youth, Later = Later In Life, BF = Blindfolded.
ID Age Gender Visual Status Onset Education Occupation AI Usage TSR (C) TSR (O) Satisfaction Scores
U1 90 Male B Youth Primary school Retired / home living No prior; needs guided onboarding 60.00% 75.00% 7 (Generally satisfied)
U2 35 Male B Youth Vocational secondary Manual work Screen reader only; willing to adopt 100.00% 66.67% 8 (Generally satisfied)
U3 30 Female SVI Youth Vocational secondary Massage therapist Limited; object/scene for tools/traffic 75.00% 50.00% 9 (Highly satisfied)
U4 21 Female SVI Youth University (undergrad) Student Occasional; study assistance 80.00% 83.33% 7 (Generally satisfied)
U5 27 Female B Later Junior college (associate) Family-assisted living Yes; object/scene, family-assisted 80.00% 100.00% 6 (Generally satisfied)
U6 24 Female S BF University Student Regular chat with study assistance 100.00% 100.00% 7 (Generally satisfied)
U7 22 Female S BF University Student Regular chat; daily queries 100.00% 100.00% 8 (Generally satisfied)
U8 24 Female S BF University Student Regular; reading augmentation 100.00% 80.00% 8 (Generally satisfied)